Interior and exterior walls of a rectangular 80,000 square foot building cost 90$ per running foot. The building is to be divided into 10 rooms by four interior walls in the x direction and one interior wall in the y direction. What should be the building dimensions if the wall cost is to be minimum.


1
Expert's answer
2021-12-15T17:24:25-0500

Let "x" be the width (in feet) and "y" be the length. Then we have

"xy=80000"

Now,counting interior and exterior walls,there are 12 walls in the "x-direction" and 6 in the "y-direction" .the number of linear feet of the wall is

"12x+6y"

We need to minimize the cost of the walls,which means minimizing the number of linear feet of the walls

First, let's express the number of linear feet of walls as a function of single variable as follows;


"y={80000\over x}"


"L(x)=12x+6.({80000\over x})=12x+{480000\over x}"

Secondly, let's find where the derivative of the function is equal to zero

"{dL\over dx}=12-{480000\over x^2}=0"

"12={480000\over x^2}"

"12x^2=480000"

"x^2=40000"

"x=200"

The total number of linear feet,and therefore the total cost of the walls, is minimum when then width is

"x=200" feet and the length is "{80000\over x}={80000\over 200}=400" feet